OpenStudy (anonymous):

Find the area of a right triangle with leg 45 and a hypotenuse of 51.

OpenStudy (anonymous):

if i remember right the equation for the area of a triangle is (base*height)/2, so use a^2 + b^2 = c^2 find the other leg and solve

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\] \[a ^{2}+45^{2}=51^{2}\] \[a ^{2}+2025=2601\] \[a ^{2}=576\] \[\sqrt{a ^{2}}=\sqrt{576}\] \[a=24\] The third leg is 24.